OpenStudy (anonymous):

WILL FAN AND MEDAL, PLEASE HELP!!!!!!!!!)= if f(x)=x-6 and g(x)=x^1/2 (x+3), find g(x) x f(x)

OpenStudy (anonymous):

is that \[g(x)\times f(x)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[Answers: ^5\sqrt{x}^2-3^2\sqrt{^2}-18x^1/2 \]

OpenStudy (anonymous):

\[\sqrt{^5}-3\sqrt{x^3}-18\sqrt{x}\]

OpenStudy (anonymous):

\[\sqrt{x^5}-\sqrt{x^3}-6\sqrt{x}\]

OpenStudy (anonymous):

\[\sqrt{x^5}-3x^3-18\sqrt{x}\]

OpenStudy (anonymous):

yea it right

OpenStudy (anonymous):

wait whats the answer?? lol